OpenStudy (anonymous):

What are the missing y-values in the table below? y = x - 2 x y 3 ? 6 ? 8 ? 10 ?

OpenStudy (anonymous):

sosssss

OpenStudy (anonymous):

all you have to do is put values of x in y= x-2 and find the value of y so when x= 3 y= 3-2 =1 now find for the rest

OpenStudy (anonymous):

y = x - 2 y = 3 - 2 y = 1 y = 6 - 2 y = 4 y = 8 - 2 y = 6 y = 10 - 2 y = 8

OpenStudy (anonymous):

thank uuuuu

OpenStudy (anonymous):

No problem :D